The cost to install new carpet in an office is determined by a $800 fixed fee plus a fee of $70 per square yard of floor space to be covered. How many square yards of floor space can be carpeted at a cost of $2500? Round to the nearest square yard.
= sq yd

Answer by fractalier(6550)

Already solved this one twice...
2500 = 800 + 70x
1700 = 70x
x = 1700 / 70 = 24.3 or about 24 square yards...